# Evaluation Guidelines for Empirical Studies in Software Engineering involving LLMs

### v1

@@ Abstract.
Large language models (LLMs) are increasingly being integrated into software engineering (SE) research and practice, yet their non-determinism, opaque training data, and evolving architectures complicate the reproduction and replication of empirical studies. We present a community effort to scope this space, introducing a taxonomy of LLM-based study types together with eight guidelines for designing and reporting empirical studies involving LLMs. The guidelines present essential (must) criteria as well as desired (should) criteria and target transparency throughout the research process. Our recommendations, contextualized by our study types, are: (1) to declare LLM usage and role; (2) to report model versions, configurations, and fine-tuning; (3) to document tool architectures; (4) to disclose prompts and interaction logs; (5) to use human validation; (6) to employ an open LLM as a baseline; (7) to report suitable baselines, benchmarks, and metrics; and (8) to openly articulate limitations and mitigations. Our goal is to enable reproducibility and replicability despite LLM-specific barriers to open science. We maintain the study types and guidelines online as a living resource for the community to use and shape (llm-guidelines.org).
@@ 1. Introduction
The importance of open science practices and documenting the study setup is not to be underestimated (Gundersen et al., 2024). One of the primary risks in creating irreproducible and irreplicable results based on studies involving LLMs stems from the variability in model performance due to their inherent non-determinism, but also due to differences in configuration, training data, model architecture, or evaluation metrics. Slight changes can lead to significantly different results. The lack of standardized benchmarks and evaluation protocols further hinders the reproducibility and replicability of study results. Without detailed information on the exact study setup, benchmarks, and metrics used, it is challenging for other researchers to replicate results using different models and tools. These issues highlight the need for clear guidelines and best practices for designing and reporting studies involving LLMs. While the SE research community has developed guidelines for conducting and reporting specific types of empirical studies such as controlled experiments (e.g., Experimentation in Software Engineering (Wohlin et al., 2024), Guide to Advanced Empirical Software Engineering (Shull et al., 2008)) or their replications (e.g., A Procedure and Guidelines for Analyzing Groups of Software Engineering Replications (Santos et al., 2021)), we believe that LLMs have specific intrinsic characteristics that require specific guidelines for researchers to achieve an acceptable level of reproducibility and replicability (see also our previous position paper (Wagner et al., 2025)). For example, even if we knew the specific version of a commercial LLM used for an empirical study, the reported task performance could still change over time, since commercial models are known to evolve beyond version identifiers (Chen et al., 2023). Moreover, commercial providers do not guarantee the availability of old model or tool versions indefinitely. In addition to version differences, LLM performance varies widely depending on configured parameters such as temperature. Therefore, not reporting the parameter settings severely impacts reproducibility. Even for “open” models such as Llama, we do not know how they were fine-tuned for specific tasks and what the exact training data was (Gibney, 2024). A general problem when evaluating LLMs’ performance is that we do not know whether the solution to a certain problem was part of the training data or not.
@@
So far, there are no holistic guidelines for conducting and reporting studies involving LLMs in SE research. With this community effort, we try to fill this gap. After outlining our Scope, we continue by introducing a taxonomy of Study Types before presenting eight Guidelines that the authors of this article co-developed. The most recent version is always available online (llm-guidelines.org); other researchers can suggest changes via a public GitHub repository.
@@ 2. Scope
First, we want to clarify that our focus is on LLMs, that is, natural language use cases. Multi-modal foundational models are beyond the scope of our study types and guidelines. We are aware that these foundational models have great potential to support software engineering research and practice. However, due to the diversity of artifacts that can be generated or used as input (e.g., images, audio, and video) and the more demanding hardware requirements, we deliberately focus on LLMs only. However, our guidelines could be extended in the future to include foundational models beyond natural language text.
@@
Second, given the exponential growth in LLM usage across all research domains, we also want to define the research contexts in which our guidelines apply. LLMs are already widely used to support several aspects of the overall research process, from fairly simple tasks such as proof-reading, spell-checking, and text translation, to more complex activities such as data coding and synthesis of literature reviews. The Study Types and Guidelines we describe are tailored to software engineering (SE) research, but we expect many of our study types to generalize beyond that domain. On the practical side, we focus on AI for software engineering (AI4SE), that is, studying the support and automation of SE tasks with the help of artificial intelligence (AI), more specifically LLMs (see Section LLMs as Tools for Software Engineers). In terms of research support, we focus on empirical SE research supported by LLMs (see Section LLMs as Tools for Software Engineering Researchers). By research support, we mean the active involvement of LLMs in data collection, processing, or analysis. We consider LLMs supporting the study design or the writing process to be out of scope.
@@
Third, our guidelines mainly target researchers planning, designing, or conducting empirical studies involving LLMs. Although researchers who review scientific articles written by others can also use our guidelines, for example, to check whether the authors adhere to the essential must requirements, reviewers are not our main target audience.
@@ 3. Study Types
The development of empirical guidelines for software engineering (SE) studies involving LLMs is crucial to ensuring the validity and reproducibility of results. However, such guidelines must be tailored to different study types that each pose unique challenges. Therefore, we developed a taxonomy of study types that we then use to contextualize the recommendations we provide. Note that our Guidelines refer to the Study Types but not the other way around.
@@ 3.1.2. LLMs as Judges
Lubos et al. evaluated the LLM’s output against human judges, to assess how well the LLM matched experts. Agreement can be measured in many ways; this study used Cohen’s kappa (Cohen, 1960) and found moderate agreement for simple requirements and poor agreement for more complex requirements. However, human evaluation of machine judgments may not be the same as evaluation of human judgments and is an open area of research (Elangovan et al., 2024). A crucial decision in studies focusing on LLMs as judges is the number of examples provided to the LLM. This might involve no tuning (zero-shot), several examples (few-shot), or providing many examples (closer to traditional training data). In the example above, Lubos et al. chose zero-shot tuning, providing no specific guidance besides the provided context; they did not show the LLM what a ‘yes’ or ‘no’ answer might look like. In contrast, Wang et al. (Wang et al., 2025) provided a rubric to an LLM when treating LLMs as a judge to produce interpretable scales (0 to 4) for the overall quality of acceptance criteria generated from user stories. They also composed this with a lower-level reward stage to refine acceptance criteria, which significantly improved the correctness, clarity, and alignment with the user stories.
@@ 3.1.3. LLMs for Synthesis
LLMs can support synthesis tasks in software engineering research by processing and distilling information from qualitative data sources. In this context, synthesis refers to the process of integrating and interpreting information from multiple sources to generate higher-level insights, identify patterns across datasets, and develop conceptual frameworks or theories. Unlike annotation (see Section LLMs as Annotators), which focuses on categorizing or labeling individual data points, synthesis involves connecting and interpreting these annotations to develop a cohesive understanding of the phenomenon being studied. Synthesis tasks can also mean generating synthetic datasets (e.g., source code, bug-fix pairs, requirements, etc.) that are then used in downstream tasks to train, fine-tune, or evaluate existing models or tools. In this case, the synthesis is done primarily based on the training data of the model; the input that researchers provide is limited to prompts and examples.
@@ 3.1.4. LLMs as Subjects
LLMs can be used as subjects in various types of empirical studies, enabling researchers to simulate human participants. The broader applicability of LLM-based studies beyond software engineering has been compiled by Xu et al. (Xu et al., 2024b), who examined various uses in social science research. Given the socio-technical nature of software development, some of these approaches are transferable to empirical software engineering research. For example, LLMs can be applied in survey and interview studies to impersonate developers responding to survey questionnaires or interviews, allowing researchers to test the clarity and effectiveness of survey items or to simulate responses under varying conditions, such as different levels of expertise or cultural contexts. For example, Gerosa et al. (Gerosa et al., 2024) explored persona-based interviews and multi-persona focus groups, demonstrating how LLMs can emulate human responses and behaviors while addressing ethical concerns, biases, and methodological challenges. Another example are usability studies, in which LLMs can simulate end-user feedback, providing insights into potential usability issues and offering suggestions for improvement based on predefined user personas. This aligns with the work of Bano et al. (Bano et al., 2025), who investigated biases in LLM-generated candidate profiles in SE recruitment processes. Their study, which analyzed both textual and visual inputs, revealed biases favoring male, Caucasian candidates, lighter skin tones, and slim physiques, particularly for senior roles.
@@ 3.2.1. Studying LLM Usage in Software Engineering
Khojah et al. investigated the use of ChatGPT (GPT-3.5) by professional software engineers in a week-long observational study (Khojah et al., 2024). They found that most developers do not use the code generated by ChatGPT directly but instead use the output as a guide to implement their own solutions. Azanza et al. conducted a case study that evaluated the impact of introducing LLMs on the onboarding process of new software developers (Azanza et al., 2024) (GPT-3). Their study identified potential in the use of LLMs for onboarding, as it can allow newcomers to seek information on their own, without the need to bother senior colleagues. Surveys can help researchers to quickly provide an overview of the current perceptions of LLM usage. For example, Jahic and Sami surveyed participants from 15 software companies regarding their practices on LLMs in software engineering (Jahic and Sami, 2024). They found that the majority of study participants had already adopted AI for software engineering tasks; most of them used ChatGPT. Multiple participants cited copyright and privacy issues, as well as inconsistent or low-quality outputs, as barriers to adoption. Retrospective studies that analyze data generated while developers use LLMs can provide additional insights into human-LLM interactions. For example, researchers can employ data mining methods to build large-scale conversation datasets, such as the DevGPT dataset introduced by Xiao et al. (Xiao et al., 2024). Conversations can then be analyzed using quantitative (Rabbi et al., 2024) and qualitative (Mohamed et al., 2024) analysis methods.
@@ 3.2.2. LLMs for New Software Engineering Tools
Traditional approaches, such as static analysis, are deterministic. LLMs are not. Although the non-determinism of LLMs can be mitigated using configuration parameters and prompting strategies, this poses a major challenge. It can be challenging for researchers to evaluate the effectiveness of a tool, as minor changes in the input can lead to major differences in the performance. Since the exact training data is often not published by model vendors, a reliable assessment of tool performance for unknown data difficult. From an engineering perspective, while open models are available, the most capable ones require substantial hardware resources. Using cloud-based APIs or relying on third-party providers for hosting, while seemingly a potential solution, introduces new concerns related to data privacy and security.
@@ 3.2.3. Benchmarking LLMs for Software Engineering Tasks
Properly built benchmarks provide objective, reproducible evaluation across different tasks, enabling a comparison between different models (and versions). In addition, benchmarks built for specific SE tasks can help identify LLM weaknesses and support their optimization and fine-tuning for such tasks. They can foster open science practices by providing a common ground for sharing data (e.g., as part of the benchmark itself) and results (e.g., of models run against a benchmark). Benchmarks built using real-world data can help legitimize research results for practitioners, supporting industry-academia collaboration. However, benchmarks are subject to several challenges.
@@
Benchmark contamination, that is, the inclusion of the benchmark in the LLM training data (Ahuja et al., 2024), has recently been identified as a problem. The careful selection of samples and the creation of the corresponding input prompts is particularly important, as correlations between prompts may bias the benchmark results (Siska et al., 2024). Although LLMs might perform well on a specific benchmark such as HumanEval, they do not necessarily perform well on other benchmarks. Moreover, benchmark metrics such as perplexity or BLEU-N do not neccessarily reflect human judgment. Recently, Cao et al. (Cao et al., 2025) has proposed guidelines for the creation of LLM benchmarks related to coding tasks, grounded in a systematic survey of existing benchmarks. In this process, they highlight current shortcomings related to reliability, transparency, irreproducibility, low data quality, and inadequate validation measures. For more details on benchmarks, see Section Report Suitable Baselines, Benchmarks, and Metrics.

### v2

# Guidelines for Empirical Studies in Software Engineering involving Large Language Models

@@ Abstract.
Large language models (LLMs) are increasingly being integrated into software engineering (SE) research and practice, yet their non-determinism, opaque training data, and evolving architectures complicate the reproduction and replication of empirical studies. We present a community effort to scope this space, introducing a taxonomy of LLM-based study types together with eight guidelines for designing and reporting empirical studies involving LLMs. The guidelines present essential (must) criteria as well as desired (should) criteria and target transparency throughout the research process. Our recommendations, contextualized by our study types, are: (1) to declare LLM usage and role; (2) to report model versions, configurations, and fine-tuning; (3) to document tool architectures; (4) to disclose prompts and interaction logs; (5) to use human validation; (6) to employ an open LLM as a baseline; (7) to use suitable baselines, benchmarks, and metrics; and (8) to openly articulate limitations and mitigations. Our goal is to enable reproducibility and replicability despite LLM-specific barriers to open science. We maintain the study types and guidelines online as a living resource for the community to use and shape (llm-guidelines.org).
@@ 1. Introduction
The importance of open science practices and documenting the study setup is not to be underestimated (Gundersen et al., 2024). One of the primary risks in creating irreproducible and irreplicable results based on studies involving LLMs stems from the variability in model performance due to their inherent non-determinism, but also due to differences in configuration, training data, model architecture, or evaluation metrics. Slight changes can lead to significantly different results. The lack of standardized benchmarks and evaluation protocols further hinders the reproducibility and replicability of study results. Without detailed information on the exact study setup, benchmarks, and metrics used, it is challenging for other researchers to replicate results using different models and tools. These issues highlight the need for clear guidelines and best practices for designing and reporting studies involving LLMs. While the SE research community has developed guidelines for conducting and reporting specific types of empirical studies such as controlled experiments (e.g., Experimentation in Software Engineering (Wohlin et al., 2024), Guide to Advanced Empirical Software Engineering (Shull et al., 2008)) or their replications (e.g., A Procedure and Guidelines for Analyzing Groups of Software Engineering Replications (Santos et al., 2021)), we believe that LLMs have specific intrinsic characteristics that require specific guidelines for researchers to achieve an acceptable level of reproducibility and replicability (see also our previous position paper (Wagner et al., 2025)). For example, even if we knew the specific version of a commercial LLM used for an empirical study, the reported task performance could still change over time, since commercial models are known to evolve beyond version identifiers (Chen et al., 2023). Moreover, commercial providers do not guarantee the availability of old model or tool versions indefinitely. In addition to version differences, LLM performance varies widely depending on configured parameters such as temperature. Therefore, not reporting the parameter settings severely impacts reproducibility. Even for “open” models such as Mistral, DeepSeek or Llama, we do not know how they were fine-tuned for specific tasks and what the exact training data was (Gibney, 2024). A general problem when evaluating LLMs’ performance is that we do not know whether the solution to a certain problem was part of the training data or not.
@@
So far, there are no holistic guidelines for conducting and reporting studies involving LLMs in SE research. With this community effort, we try to fill this gap. After outlining our Scope, we continue by introducing a taxonomy of Study Types before presenting eight Guidelines that the authors of this article co-developed. Also, for each study type and guideline, we identify relevant research exemplars, both within the SE literature, but also outside the discipline where relevant; this serves a useful purpose given the vast proliferation of research on this nascent topic. The most recent version is always available online. [^1] https://llm-guidelines.org Other researchers can suggest changes via a public GitHub repository.xw [^2] https://github.com/se-uhd/llm-guidelines-website
@@ 2. Scope
In this section, we outline the scope of the study types and guidelines we present.
SE as our Target Discipline
We target software engineering (SE) research conducted in academic or industry contexts. While other disciplines have started developing community guidelines for reporting empirical studies involving LLMs, SE has not seen a holistic effort in that direction. In healthcare, for example, Gallifant et al. presented the TRIPOD-LLM guidelines (Gallifant et al., 2025), which are centered around a checklist for good reporting of studies that are developing, tuning, prompt engineering or evaluating an LLM (Gallifant et al., 2025). Although the checklist overlaps with our Guidelines (e.g., reporting the LLM name and version, comparing performance between LLMs, humans and other benchmarks), they are healthcare-specific (e.g., they require to explain the healthcare context and therapeutic and clinical workflow). Instead of a checklist, we distinguish must and should criteria and provide brief tl;dr summaries highlighting our most important recommendations. SE research utilizes a broader range of empirical methods than many other disciplines. Therefore, we are convinced that SE-specific guidelines need to be developed in combination with a Study Types taxonomy such as ours. In SE, Sallou et al. have presented a vision paper on mitigating validity threats in LLM-based SE research. Although their guidelines partially overlap with ours (e.g., repeating experiments to handle output variability, considering data leakage), our guidelines cover a broader range of study types and provide more detailed advise. For example, we contextualize our guidelines by study type, distinguish LLMs and LLM-based tools, talk about models, baselines, benchmarks, and metrics, and explicitly address human validation. Moreover, our guidelines are the result of an extensive coordination process between a large number of experts in empirical SE, hence a first step towards true community guidelines.
Focus on Natural Language Use Cases
@@
We want to clarify that our focus is on LLMs, that is, natural language use cases. Multi-modal foundational models are beyond the scope of our study types and guidelines. We are aware that these foundational models have great potential to support SE research and practice. However, due to the diversity of artifacts that can be generated or used as input (e.g., images, audio, and video) and the more demanding hardware requirements, we deliberately focus on LLMs only. However, our guidelines could be extended in the future to include foundational models beyond natural language text.
Focus on Direct Tool or Research Support
@@
Given the exponential growth in LLM usage across all research domains, we also want to define the research contexts in which our guidelines apply. LLMs are already widely used to support several aspects of the overall research process, from fairly simple tasks such as proof-reading, spell-checking, and text translation, to more complex activities such as data coding and synthesis of literature reviews. Regarding tools, we focus on use cases in the area of AI for software engineering (AI4SE), that is, studying the support and automation of SE tasks with the help of artificial intelligence (AI), more specifically LLMs (see Section LLMs as Tools for Software Engineers). For research support, we focus on empirical SE research supported by LLMs (see Section LLMs as Tools for Software Engineering Researchers). By research support, we mean the active involvement of LLMs in data collection, processing, or analysis. We consider LLMs supporting the study design or the writing process to be out of scope.
Researchers as our Target Audience
@@
Third, our guidelines mainly target SE researchers planning, designing, conducting, and reporting empirical studies involving LLMs. Although researchers who review scientific articles written by others can also use our guidelines, for example, to check whether the authors adhere to the essential must requirements we present, reviewers are not our main target audience.
3. Methodology
[…]
We held bi-weekly meetings, in which we decided on the structure of the study type taxonomy and the guidelines. During these discussions, we added a new study type (LLMs for Synthesis), refined the guidelines to Report Model Version, Configuration, and Customizations and to Report Tool Architecture beyond Models, extended the guideline to Report Prompts, their Development, and Interaction Logs, and added the new guidelines to Use Suitable Baselines, Benchmarks, and Metrics and to Report Limitations and Mitigations. Then, at least two co-authors were assigned to each study type and guideline. They refined and extended the content to align with the intended structure, followed by a peer review and refinement phase. Afterwards, the first author reviewed all sections and discussed potential changes with the topic leads. After completing the internal review, we invited external experts in empirical SE to review the study types and guidelines. After incorporating their feedback, [^3] Discussion on GitHub: https://github.com/se-uhd/llm-guidelines-website/issues/89 we compiled and published this article.
@@ 4. Study Types
The development of empirical guidelines for software engineering (SE) studies involving LLMs is crucial to ensuring the validity and reproducibility of results. However, such guidelines must be tailored to different study types that each pose unique challenges. Therefore, we developed a taxonomy of study types that we then use to contextualize the recommendations we provide.
@@ 4.1.2. LLMs as Judges
Lubos et al. evaluated the LLM’s output against human judges, to assess how well the LLM matched experts. Agreement can be measured in many ways; this study used Cohen’s kappa (Cohen, 1960) and found moderate agreement for simple requirements and poor agreement for more complex requirements. However, human evaluation of machine judgments may not be the same as evaluation of human judgments and is an open area of research (Elangovan et al., 2024). A crucial decision in studies focusing on LLMs as judges is the number of examples provided to the LLM. This might involve no tuning (zero-shot), several examples (few-shot), or providing many examples (closer to traditional training data). In the example above, Lubos et al. chose zero-shot tuning, providing no specific guidance besides the provided context; they did not show the LLM what a ‘yes’ or ‘no’ answer might look like. In contrast, Wang et al. (Wang et al., 2025) provided a rubric to an LLM when treating LLMs as a judge to produce interpretable scales (0 to 4) for the overall quality of acceptance criteria generated from user stories. They combined this with a lower-level reward stage to refine acceptance criteria, which significantly improved correctness, clarity, and alignment with the user stories.
@@ 4.1.3. LLMs for Synthesis
LLMs can support synthesis tasks in software engineering research by processing and distilling information from qualitative data sources. In this context, synthesis refers to the process of integrating and interpreting information from multiple sources to generate higher-level insights, identify patterns across datasets, and develop conceptual frameworks or theories. Unlike annotation (see Section LLMs as Annotators), which focuses on categorizing or labeling individual data points, synthesis involves connecting and interpreting these annotations to develop a cohesive understanding of the phenomenon being studied. Synthesis tasks can also mean generating synthetic datasets (e.g., source code, bug-fix pairs, requirements, etc.) that are then used in downstream tasks to train, fine-tune, or evaluate existing models or tools. In this case, the synthesis is done primarily using the LLM and its training data; the input is limited to basic instructions and examples.
@@ 4.1.4. LLMs as Subjects
LLMs can be used as subjects in various types of empirical studies, enabling researchers to simulate human participants. The broader applicability of LLM-based studies beyond software engineering has been compiled by Xu et al. (Xu et al., 2024b), who examined various uses in social science research. Given the socio-technical nature of software development, some of these approaches are transferable to empirical software engineering research. For example, LLMs can be applied in survey and interview studies to impersonate developers responding to survey questionnaires or interviews, allowing researchers to test the clarity and effectiveness of survey items or to simulate responses under varying conditions, such as different levels of expertise or cultural contexts. For example, Gerosa et al. (Gerosa et al., 2024) explored persona-based interviews and multi-persona focus groups, demonstrating how LLMs can emulate human responses and behaviors while addressing ethical concerns, biases, and methodological challenges. Another example is usability studies, in which LLMs can simulate end-user feedback, providing insights into potential usability issues and offering suggestions for improvement based on predefined user personas. This aligns with the work of Bano et al. (Bano et al., 2025), who investigated biases in LLM-generated candidate profiles in SE recruitment processes. Their study, which analyzed both textual and visual inputs, revealed biases favoring male Caucasian candidates, lighter skin tones, and slim physiques, particularly for senior roles.
@@ 4.2.1. Studying LLM Usage in Software Engineering
Based on a convergent mixed-methods study, Russo has found that early adoption of generative AI by software engineers is primarily driven by compatibility with existing workflows (Russo, 2024). Khojah et al. investigated the use of ChatGPT (GPT-3.5) by professional software engineers in a week-long observational study (Khojah et al., 2024). They found that most developers do not use the code generated by ChatGPT directly but instead use the output as a guide to implement their own solutions. Azanza et al. conducted a case study that evaluated the impact of introducing LLMs on the onboarding process of new software developers (Azanza et al., 2024) (GPT-3). Their study identified potential in the use of LLMs for onboarding, as it can allow newcomers to seek information on their own, without the need to bother senior colleagues. Jahic and Sami surveyed participants from 15 software companies regarding their practices on LLMs in software engineering (Jahic and Sami, 2024). They found that the majority of study participants had already adopted AI for software engineering tasks; most of them used ChatGPT. Multiple participants cited copyright and privacy issues, as well as inconsistent or low-quality outputs, as barriers to adoption. Retrospective studies that analyze data generated while developers use LLMs can provide additional insights into human-LLM interactions. For example, researchers can employ data mining methods to build large-scale conversation datasets, such as the DevGPT dataset introduced by Xiao et al. (Xiao et al., 2024). Conversations can then be analyzed using quantitative (Rabbi et al., 2024) and qualitative (Mohamed et al., 2024) analysis methods.
@@ 4.2.2. LLMs for New Software Engineering Tools
Traditional approaches, such as static analysis, are deterministic. LLMs are not. Although the non-determinism of LLMs can be mitigated using configuration parameters and prompting strategies, this poses a major challenge. It can be challenging for researchers to evaluate the effectiveness of a tool, as minor changes in the input can lead to major differences in the performance. Since the exact training data is often not published by model vendors, a reliable assessment of tool performance for unknown data is difficult. From an engineering perspective, while open models are available, the most capable ones require substantial hardware resources. Using cloud-based APIs or relying on third-party providers for hosting, while seemingly a potential solution, introduces new concerns related to data privacy and security.
@@ 4.2.3. Benchmarking LLMs for Software Engineering Tasks
Properly built benchmarks provide objective, reproducible evaluation across different tasks, enabling a comparison between different models (and versions). In addition, benchmarks built for specific SE tasks can help identify LLM weaknesses and support their optimization and fine-tuning for such tasks. They can foster open science practices by providing a common ground for sharing data (e.g., as part of the benchmark itself) and results (e.g., of models run against a benchmark). Benchmarks built using real-world data can help legitimize research results for practitioners, supporting industry-academia collaboration.
@@
Benchmark contamination, that is, the inclusion of the benchmark in the LLM training data (Ahuja et al., 2024), has recently been identified as a problem. The careful selection of samples and the creation of the corresponding input prompts is particularly important, as correlations between prompts may bias the benchmark results (Siska et al., 2024). Although LLMs might perform well on a specific benchmark such as HumanEval or SWE-bench, they do not necessarily perform well on other benchmarks. Moreover, benchmarks usually do not capture the full complexity and diversity of software engineering work (Chandra, 2025). Another issue is that metrics used for benchmarks, such as perplexity or BLEU-N, do not necessarily reflect human judgment. Recently, Cao et al. (Cao et al., 2025) have proposed guidelines for the creation of LLM benchmarks related to coding tasks, grounded in a systematic survey of existing benchmarks. In this process, they highlight current shortcomings related to reliability, transparency, irreproducibility, low data quality, and inadequate validation measures. For more details on benchmarks, see Section Use Suitable Baselines, Benchmarks, and Metrics.